# Ab initio uncertainty quantification in scattering analysis of microscopy

### v2

@@ III.1 A latent factor model of isotropic processes
We initially direct our attention to isotropic processes, where the ISF is the same at each ring of pixels in the reciprocal space. Extension to anisotropic processes will be discussed in Section III.5. Consider a latent factor model of an $N=N_{1}\times N_{2}$ pixels of real-valued image intensity $\mathbf{y}(t)=[y(\mathbf{x}_{1},t),...,y(\mathbf{x}_{N},t)]^{T}$ at time $t$ below
@@
where $\bm{\epsilon}(t)\sim\mathcal{MN}(\mathbf{0},\frac{\bar{B}}{2}\mathbf{I}_{N})$ is an N-dimensional Gaussian white noise vector with variance $\frac{\bar{B}}{2}$ and $\mathbf{I}_{N}$ being the identity matrix of N dimensions, the $N\times N$ matrix $\mathbf{W}^{*}$ is a 2D inverse Fourier basis (or complex conjugate of the Fourier basis), which relates the $N$ observations of an image at time $t$ from Cartesian space $\mathbf{x}=(x_{1},x_{2})^{T}$ to a set of random factor processes $\mathbf{z}(t)$ in the reciprocal space $\mathbf{q}=(q_{1},q_{2})^{T}$. The latent factor $\mathbf{z}(t)$ is an $N$ dimensional complex random vector: $\mathbf{z}(t)=\mathbf{z}_{re}(t)+i\mathbf{z}_{im}(t)$, with each random factor at $n$ time points independently following a zero-mean multivariate normal distribution: $\mathbf{z}_{j^{\prime},re}\sim\mathcal{MN}(\mathbf{0},\frac{A_{j}}{4}\mathbf{R}_{j})$ and $\mathbf{z}_{j^{\prime},im}\sim\mathcal{MN}(\mathbf{0},\frac{A_{j}}{4}\mathbf{R}_{j})$ for $j=1,...,J$. The $(k_{1},k_{2})$th entry of $\mathbf{R}_{j}$ is characterized by ISF: $R_{j}(k_{1},k_{2})=f_{\bm{\theta}}(q_{j},\Delta t_{k})$ with $\Delta t_{k}=|k_{2}-k_{1}|\Delta t_{min}$ with $\Delta t_{min}$ being the interval between two consecutive time frames. The key is that the correlation matrix $\mathbf{R}_{j}$ of the latent factor is formed by the ISF from the physical process. This means each entry of the real and imaginary random factors corresponds to one Fourier-transformed quantity, where the covariance is parameterized by the amplitude and intermediate scattering function: $\mathbb{E}[z_{\mathbf{q},re}(t)z_{\mathbf{q},re}(t+\Delta t)]=\mathbb{E}[z_{\mathbf{q},im}(t)z_{\mathbf{q},im}(t+\Delta t)]=\frac{A(\mathbf{q})}{4}f_{\bm{\theta}}(\mathbf{q},\Delta t)$. Without loss of generality, we assume that $N_{1}=N_{2}=\sqrt{N}$, i.e. square image at each time frame.
@@ III.2 DDM is fitting the temporal variogram of the latent factor model in the reciprocal space
which is the mean of the image structure function in Eq. (3), the statistics used for estimating parameters in DDM. Eq. (8) means that if we assume the generative model in Eq. (5), the expected value of the image structure function in Eq. (3) is the expected value of ${({\hat{y}}_{\mathbf{q}}(t+\Delta t)-{\hat{y}}_{\mathbf{q}}(t))({\hat{y}}^{*}_{\mathbf{q}}(t+\Delta t)-{\hat{y}}^{*}_{\mathbf{q}}(t))}$, equivalent to a temporal variagram in the reciprocal space. Fitting a spatial variogram in the real space was extensively studied by the statistics community [46, 47, 48]. However, it can be strenuous to fit the temporal variogram at each wavevector $\mathbf{q}$ in the reciprocal space and optimally aggregate the estimators, since the variogram is correlated at each lag time and the amplitude parameter is drastically different at distinct $\mathbf{q}$. Thus, directly fitting the temporal variogram in the reciprocal space and aggregating the estimators could lead to unstable estimation. Next, we will introduce the maximum marginal likelihood estimator of the parameters, which provides a natural and optimal way to aggregate information on each wavevector.
@@ III.3 Maximum marginal likelihood estimator
where $\bm{\Sigma}_{j}=\frac{A_{j}}{4}\mathbf{R}_{j}+\frac{\bar{B}}{4}\mathbf{I}_{n}$. Here $\mathcal{S}_{j}$ denotes the index set of the $j$th ring of isotropic processes, for $j=1,...,J$, and $p_{MN}(\mathbf{s};\bm{\mu},\bm{\Sigma})$ denotes the density of an n-vector multivariate normal distribution at real values $\mathbf{s}$ with mean and covariance being $\bm{\mu}$ and $\bm{\Sigma}$ respectively: $p_{MN}(\mathbf{s};\bm{\mu},\bm{\Sigma})=(2\pi)^{-n/2}|\bm{\Sigma}|^{-1/2}\exp\left\{-\frac{1}{2}(\mathbf{s}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{s}-\bm{\mu})\right\}$. The derivation of Eq. (9) is given in Appendix C.
We denote $S_{j}=\#\mathcal{S}_{j}$, the number of transformed pixels within the index set $\mathcal{S}_{j}$, and let the total number of pixels within $J$ rings be $\tilde{N}=\sum^{J}_{j=1}S_{j}$. Note that we do not include the transformed output outside of the $J$ rings in the likelihood function, consistent with DDM, and hence $\tilde{N}<N$. In principle, one can compute the likelihood of all quantities inside or outside the $J$ rings. However, as the transformed quantities at large wavevectors become similar to noise, we may only need to include a small number of wavevectors, which is much smaller than $J$. Based on this feature, we will introduce feasible ways to further reduce the computation and storage cost in Section IV.2. Using all wavevectors for estimation is implemented as the default version of our software packages [32, 33]; a reduction of the wavevector range can be executed to decrease computational cost.
@@
$A_{est,j}=\frac{2}{S_{j}n}\sum_{j^{\prime}\in\mathcal{S}_{j}}\sum^{n}_{k=1}|\hat{y}_{j^{\prime}}(t_{k})|^{2}-{\bar{B}}$ (10)
for $j=1,...,J$. In practice, we may take the absolute value of $A_{est,j}$ for some of the large $j$ to keep it nonnegative. Then we estimate the model parameter $\bm{\theta}$ and noise parameter $\bar{B}$ by maximizing the marginal likelihood with the estimated $\mathbf{A}_{est,1:J}$ from Eq. (10)
$(\bm{\theta}_{est},\bar{B}_{est})=\mbox{argmax}_{\bm{\theta},\bar{B}}\mathcal{L}\left(\bm{\theta},\mathbf{A}_{est,1:J},\bar{B}\right).$ (11)
This step dramatically reduces the dimension of the parameter space and makes numerical optimization much more stable than optimizing in a large parameter space. We use the low-storage quasi-Newton optimization method [49] for estimation $\bm{\theta}$ and $\bar{B}$ by maximizing the logarithm of the marginal likelihood function $\mbox{log}(\mathcal{L}\left(\bm{\theta},\mathbf{A}_{est,1:J},\bar{B}\right))$ in Eq. (9).
Eq. (5) defines a generative model for the untransformed data from the beginning of the analysis, and the latent factor processes are integrated out to propagate the uncertainty to derive the marginal likelihood in Eq. (9) for parameter estimation. Thus we call the method ab initio uncertainty quantification (AIUQ) for scattering analysis of microscopy. Note here we do not need to compute the difference of image pairs as in the image structure function in Eq. (3); instead, we will apply the generalized Schur algorithm [50, 51, 52] for accelerating the computation of Toeplitz covariances from the marginal likelihood in Eq. (9) summarized in Section IV.1 and Section S1 in SI.
@@
Figure 1: (a) B, D estimates in likelihood space using different methods: AIUQ (x), AIUQ with all q (o), $D(q,\Delta t)$ fit 1 ($\triangle$), $D(q,\Delta t)$ fit 2 ($\diamond$), whereas the solid dot denotes the truth. The color is the log marginal likelihood by the AIUQ approach with reduced $q$. (b) The estimate of the diffusion coefficient $\theta$ using different methods with the fraction of $q$ used in the estimate.
@@
We illustrate the estimation accuracy of AIUQ using a slow diffusive process in Fig. 1. When dealing with slow dynamics, where the standard deviation of the displacement is smaller than 1/10 of a pixel in each timestep, MPT sometimes fails to accurately capture the underlying dynamics at small lag times, due to the similarity between noise and signal. We compare AIUQ with two DDM approaches. In the first DDM approach, termed $D(q,\Delta t)$ fit 1, the noise parameter is estimated by $\bar{B}_{est}=D(\mathbf{q}_{min},\Delta t_{min})$, and the amplitude parameter $A_{j}$ is estimated by the unbiased estimator specified in Eq. (10). Then these parameters are substituted into Eq. (4), so that the parameters in ISF are numerically optimized by minimizing the $L_{2}$ loss between the reconstructed and observed image structure functions. In the second DDM approach, termed $D(q,\Delta t)$ fit 2, parameters are separately estimated for wavevectors by numerically minimizing the $L_{2}$ loss function of the image structure function in Eq. (4). Then the average of the estimation is used to estimate the parameters. In both approaches, we find that the estimation of noise parameter $\bar{B}$ is reasonably good (Fig 1 (a)). Furthermore, $\bar{B}$ is also reasonably good compared with the truth in a wide range of simulation studies shown in Table S3 in SI. However, the estimation of diffusion coefficient by two DDM approaches is not satisfying when the entire wavevector range is used in estimation. AIUQ approaches, with all wavevectors or with a reduced set of wavevectors, use the first $J_{0}$ sets that explain $99\%$ of the variability: $({\sum^{J_{0}}_{j=1}A_{j}})/({\sum^{J}_{j=1}A_{j}})\geq 99\%$, are compared. As demonstrated in Fig. 1(a), AIUQ is more accurate in estimating the diffusion coefficient than two approaches that fit the image structure function, while performing similarly well in estimating the noise parameter. Fig. 1(b) shows fitting the image structure function in DDM can lead to q-dependent estimation. We further show another example when we directly invert the image structure function in Fig. 9 for estimation, which demonstrated the q-dependence for estimation. In comparison, the estimated parameter is not dependent on $q$ using AIUQ, when a sufficient number of wavevectors is used, shown in Fig. 1(b), as the information is naturally weighed by the marginal likelihood in Eq. (9).
@@ IV.1 The generalized Schur method of accelerating computation for Toeplitz covariances
Directly computing the marginal likelihood in Eq. (9) is slow as computing each of $n$-vector multivariate normal requires $\mathcal{O}(n^{3})$ computational operations, from computing the matrix inversion and determinant of the covariance matrix. Assuming we have $J$ rings of Fourier-transformed intensity, and the number of indices in each ring is $S_{j}$, the total computational operations scales as $\mathcal{O}(Jn^{3})+\mathcal{O}(\tilde{N}n^{2})$ operations for isotropic processes, and $\mathcal{O}(\tilde{N}n^{3})$ for anisotropic processes. As the likelihood function needs to be computed tens of times for numerically optimizing the parameters, the computational operations can be up to $10^{15}$ for a regular video with $500\times 500$ pixels and $500$ time frames, which is too computationally expensive. Here, we introduce a fast approach that can substantially reduce the computational cost without any approximation to the likelihood function.
@@ IV.2 Data reduction
For processes with an increased MSD with respect to the increase of lag time, such as subdiffusion or superdiffusion, the transformed intensity rapidly decorrelates at large wavevectors, making it indistinguishable from noise. A typical way for dimension reduction is to choose the first $J_{0}$ rings of Fourier-transformed intensity that explains the most variability of the data: $({\sum^{J_{0}}_{j=1}A_{j}})/({\sum^{J}_{j=1}A_{j}})\geq 1-\varepsilon$, where $\varepsilon$ is a small number, similar to the probabilistic principal component analysis [63]. For instance, choosing $\varepsilon=0.001$ means we have $99.9\%$ of the variability from the transformed intensity explained. As $A_{j}$ becomes close to 00 at large $j$ and the ‘ring’ gets larger for large $j$, such a choice can substantially reduce the storage and computational requirements by avoiding computing a large number of transformed intensities at high-frequency wavevectors. Another way is to ensure that we select a relatively large proportion ($J_{0}/J\geq\beta$) to retain enough information from the signal. For some confined processes, such as the Ornstein–Uhlenbeck (OU) process, where the MSD approaches a plateau at large lag time, using a large fraction of wavevectors is a safer choice, as the temporal correlation of transformed intensities does not decrease at large wavevectors. Having a more conservative threshold that selects a larger $J_{0}$, e.g. $\varepsilon=0.001$ and $\beta=0.5$, is also better for estimating the noise parameter. Here, we emphasize that the selection of $J_{0}$ is mainly for further reducing the storage and computational cost, which is different from selecting the range of wavevectors to analyze in minimizing the loss function in Eq. (4) in DDM.

### v4

@@ III.1 A latent factor model of isotropic processes
Let us first direct our attention to isotropic processes, where the ISF is the same at each ring of pixels in the reciprocal space. Extension to anisotropic processes will be discussed in Section III.5. Consider a latent factor model of an $N=N_{1}\times N_{2}$ pixels of real-valued image intensity $\mathbf{y}(t)=[y(\mathbf{x}_{1},t),...,y(\mathbf{x}_{N},t)]^{T}$ at time $t$ below
@@
where $\bm{\epsilon}(t)\sim\mathcal{MN}(\mathbf{0},\frac{\bar{B}}{2}\mathbf{I}_{N})$ is an N-dimensional Gaussian white noise vector with variance $\frac{\bar{B}}{2}$ and $\mathbf{I}_{N}$ being the identity matrix of N dimensions, the $N\times N$ matrix $\mathbf{W}^{*}$ is a 2D inverse Fourier basis (or complex conjugate of the Fourier basis), which relates the $N$ observations of an image at time $t$ from Cartesian space $\mathbf{x}=(x_{1},x_{2})^{T}$ to a set of random factor processes $\mathbf{z}(t)$ in the reciprocal space $\mathbf{q}=(q_{1},q_{2})^{T}$. The latent factor $\mathbf{z}(t)$ is an $N$ dimensional complex random vector: $\mathbf{z}(t)=\mathbf{z}_{re}(t)+i\mathbf{z}_{im}(t)$, with each random factor at $n$ time points independently following a zero-mean multivariate normal distribution: $\mathbf{z}_{\mathbf{j}^{\prime},re}\sim\mathcal{MN}(\mathbf{0},\frac{A_{j}}{4}\mathbf{R}_{j})$ and $\mathbf{z}_{\mathbf{j}^{\prime},im}\sim\mathcal{MN}(\mathbf{0},\frac{A_{j}}{4}\mathbf{R}_{j})$ for $j=1,...,J$ and for any index $\mathbf{j}^{\prime}\in\mathcal{S}_{j}$. The $(k_{1},k_{2})$th entry of $\mathbf{R}_{j}$ is characterized by ISF: $R_{j}(k_{1},k_{2})=f_{\bm{\theta}}(q_{j},\Delta t_{k})$ with $\Delta t_{k}=|k_{2}-k_{1}|\Delta t_{min}$ with $\Delta t_{min}$ being the interval between two consecutive time frames. The correlation matrix $\mathbf{R}_{j}$ encodes the two-time correlation function in the Fourier space [47], with the $(k_{1},k_{2})$ term of the covariance being $\frac{A_{j}}{4}R_{j}(k_{1},k_{2})=\mbox{Cov}\left({\hat{y}}_{q_{\mathbf{j}^{\prime}}}(t_{k_{1}}),{\hat{y}}_{q_{\mathbf{j}^{\prime}}}(t_{k_{2}})\right)$ and for a time-invariant process, the quantity reduces to the image correlation function with $\Delta t=|t_{k_{2}}-t_{k_{1}}|$ in [24]. The key is that the correlation matrix $\mathbf{R}_{j}$ of the latent factor is formed by the ISF from the physical process. This means each entry of the real and imaginary random factors corresponds to one Fourier-transformed quantity, where the covariance is parameterized by the amplitude and intermediate scattering function: $\mathbb{E}[z_{\mathbf{q},re}(t)z_{\mathbf{q},re}(t+\Delta t)]=\mathbb{E}[z_{\mathbf{q},im}(t)z_{\mathbf{q},im}(t+\Delta t)]=\frac{A(\mathbf{q})}{4}f_{\bm{\theta}}(\mathbf{q},\Delta t)$. Without loss of generality, we assume that $N_{1}=N_{2}=\sqrt{N}$, i.e. square image at each time frame.
@@ III.2 DDM is fitting the temporal variogram of the latent factor model in the reciprocal space
which is the mean of the image structure function in Eq. (3), the statistics used for estimating parameters in DDM. Eq. (8) means that if we assume the probabilistic model in Eq. (5), the expected value of the image structure function in Eq. (3) is the expected value of ${({\hat{y}}_{\mathbf{q}}(t+\Delta t)-{\hat{y}}_{\mathbf{q}}(t))({\hat{y}}^{*}_{\mathbf{q}}(t+\Delta t)-{\hat{y}}^{*}_{\mathbf{q}}(t))}$, equivalent to a temporal variagram in the reciprocal space. Fitting a spatial variogram in the real space was extensively studied by the statistics community [48, 49, 50]. However, it can be strenuous to fit the temporal variogram at each wave vector $\mathbf{q}$ in the reciprocal space and optimally aggregate the estimators, since the variogram is correlated at each lag time and the amplitude parameter can be drastically different at distinct $\mathbf{q}$. Thus, directly fitting the temporal variogram in the reciprocal space and aggregating the estimators could lead to unstable estimation. Next, we will introduce the maximum marginal likelihood estimator of the parameters, which provides a natural and optimal way to aggregate information on each wave vector.
@@ III.3 Maximum marginal likelihood estimator
where $\bm{\Sigma}_{j}=\frac{A_{j}}{4}\mathbf{R}_{j}+\frac{\bar{B}}{4}\mathbf{I}_{n}$. Here $\mathcal{S}_{j}$ denotes the index set of the $j$th ring of isotropic processes, for $j=1,...,J$, and $p_{MN}(\mathbf{s};\bm{\mu},\bm{\Sigma})$ denotes the density of an n-vector multivariate normal distribution at a real-valued vector $\mathbf{s}$ with mean and covariance being $\bm{\mu}$ and $\bm{\Sigma}$ respectively:
$p_{MN}(\mathbf{s};\bm{\mu},\bm{\Sigma})$ $=(2\pi)^{-\frac{n}{2}}|\bm{\Sigma}|^{-\frac{1}{2}}\times$
$\qquad\exp\left\{-\frac{1}{2}(\mathbf{s}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{s}-\bm{\mu})\right\}.$
The derivation of Eq. (9) is given in Appendix C.
We denote $S_{j}=\#\mathcal{S}_{j}$, the number of transformed pixels within the index set $\mathcal{S}_{j}$, and let the total number of pixels within $J$ rings be $\tilde{N}=\sum^{J}_{j=1}S_{j}$. Note that we do not include the transformed output outside of the $J$ rings in the likelihood function, consistent with DDM, and hence $\tilde{N}<N$. In principle, one can compute the likelihood of all quantities inside or outside the $J$ rings. However, as the transformed quantities at large wave vectors become similar to noise, we may only need to include a small number of wave vectors, which is much smaller than $J$. Based on this feature, we will introduce feasible ways to further reduce the computation and storage cost in Section IV.2. Using all wave vectors for estimation is implemented as the default version of our software packages [33, 34]; a reduction of the wave-vector range can be executed to decrease computational cost.
@@
$A_{est,j}=\frac{2}{S_{j}n}\sum_{\mathbf{j}^{\prime}\in\mathcal{S}_{j}}\sum^{n}_{k=1}|\hat{y}_{\mathbf{j}^{\prime}}(t_{k})|^{2}-{\bar{B}}$ (10)
for $j=1,...,J$. The unbiasedness of the estimator $A_{est,j}$ is derived in Appendix C. In practice, we may take the absolute value of $A_{est,j}$ to keep it nonnegative. Another way is to set those negative $A_{est,j}$ to 0. As $A_{est,j}$ is rarely negative and the absolute value of the negative $A_{est,j}$ is typically very small, they almost have no impact on the likelihood function. Thus, these two estimations do not change the parameter estimation significantly, as shown in Table S4 in the Supplemental Material. In our package, we offer users the option to use absolute values or setting $A_{est,j}$ to 0. Then we estimate the model parameter $\bm{\theta}$ and noise parameter $\bar{B}$ by maximizing the marginal likelihood with the estimated $\mathbf{A}_{est,1:J}$ from Eq. (10)
$(\bm{\theta}_{est},\bar{B}_{est})=\mbox{argmax}_{\bm{\theta},\bar{B}}\mathcal{\mbox{missing}}{log}\left[L\left(\bm{\theta},\mathbf{A}_{est,1:J},\bar{B}\right)\right].$ (11)
The plug-in estimator $\mathbf{A}_{est,1:J}$ dramatically reduces the dimension of the parameter space and makes numerical optimization much more stable than optimizing in a large parameter space. We use the low-storage quasi-Newton optimization method [51] for estimation $\bm{\theta}$ and $\bar{B}$ by maximizing the logarithm of the marginal likelihood function $\mbox{log}\left[\mathcal{L}\left(\bm{\theta},\mathbf{A}_{est,1:J},\bar{B}\right)\right]$ in Eq. (11).
Eq. (5) defines a generative model for the untransformed data from the beginning of the analysis, and the latent factor processes are integrated out to propagate the uncertainty to derive the marginal likelihood in Eq. (9) for parameter estimation. Thus we call the method ab initio uncertainty quantification (AIUQ) for scattering analysis of microscopy. Note here we do not need to compute the difference of image pairs as in the image structure function in Eq. (3); instead, we will apply the generalized Schur algorithm [52, 53, 54] for accelerating the computation of Toeplitz covariances from the marginal likelihood in Eq. (9) summarized in Section IV.1 and Section S1 in SI. The AIUQ approach provides a global fit of the entire data at the untransformed real space and time $\{\mathbf{x},t\}$ based on the model in Eq. (5). In comparison, DDM fits the image structure function in the Fourier space in Eq. (3) at the lag time $\{\mathbf{q},\Delta t\}$, and selecting suitable wave-vector ranges may be required to obtain an accurate estimation of the ISF.
@@
Figure 1: (a) For a simulated image stack, B, $\theta$ estimates using different methods: AIUQ (x), AIUQ with all q (o), $D(q,\Delta t)$ fit 1 ($\triangle$), $D(q,\Delta t)$ fit 2 ($\diamond$), whereas the solid dot denotes the truth. The colormap represents the log marginal likelihood by the AIUQ approach with reduced $q$. (b) The estimated diffusion coefficient $\theta$ using different methods with the fraction of $q$ used in estimation. The axes are plotted on logarithmic scales.
@@
We illustrate the estimation accuracy of AIUQ using a slow diffusive process in Fig. 1. When dealing with slow dynamics, where the standard deviation of the displacement is smaller than 1/10 of a pixel in each timestep, MPT sometimes fails to accurately capture the underlying dynamics at small lag times, due to the similarity between noise and signal. We compare AIUQ with two DDM approaches both based on all wave vectors. In the first DDM approach, termed $D(q,\Delta t)$ fit 1, the noise parameter is estimated by $\bar{B}_{est}=D(\mathbf{q}_{min},\Delta t_{min})$, and the amplitude parameter $A_{j}$ is estimated by the unbiased estimator specified in Eq. (10). Then these parameters are substituted into Eq. (4), so that the parameters in ISF $\bm{\theta}$ are numerically optimized by minimizing the $L_{2}$ loss between the reconstructed and observed image structure functions by summing up the loss at all wave vectors. In the second DDM approach, termed $D(q,\Delta t)$ fit 2, parameters are separately estimated for wave vectors by numerically minimizing the $L_{2}$ loss function of the image structure function in Eq. (4) for each $q_{j}$ to obtain $\bm{\theta}(q_{j})$, and then the average of the estimation of $\bm{\theta}_{j}(q_{j})$ is used to estimate the parameters $\bm{\theta}$. In both approaches, we find that the estimation of noise parameter $\bar{B}$ is reasonably good (Fig 1 (a)). Furthermore, $\bar{B}$ is also reasonably good compared with the truth in a wide range of simulation studies shown in Table S3 in the Supplemental Material. However, the estimation of diffusion coefficient by two DDM approaches is not satisfying when the entire wave-vector range is used in estimation. AIUQ approaches, with all wave vectors or with a reduced set of wave vectors, use the first $J_{0}$ sets that explain $99\%$ of the variability: $({\sum^{J_{0}}_{j=1}A_{j}})/({\sum^{J}_{j=1}A_{j}})\geq 99\%$, are compared. As demonstrated in Fig. 1(a), AIUQ is more accurate in estimating the diffusion coefficient than two approaches that fit the image structure function, while performing similarly well in estimating the noise parameter. Fig. 1(b) shows fitting the image structure function in DDM can lead to q-dependent estimation. We further show another example when we directly invert the image structure function in Fig. 9 for estimation, which demonstrated the q-dependence for estimation. In comparison, the estimated parameter is not dependent on $q$ using AIUQ, when a sufficient number of wave vectors is used, shown in Fig. 1(b), as the information is naturally weighed by the marginal likelihood in Eq. (9). Selecting a suitable wave-vector range for estimation can improve the estimation by DDM, whereas a principled way of selecting such range was not found.
@@ IV.1 The generalized Schur method of accelerating computation for Toeplitz covariances
Directly computing the marginal likelihood in Eq. (9) is slow as computing each of $n$-vector multivariate normal density requires $\mathcal{O}(n^{3})$ computational operations, from computing the matrix inversion and determinant of the covariance matrix. Assuming we have $J$ rings of Fourier-transformed intensity, and the number of indices in each ring is $S_{j}$, the total computational operations scales as $\mathcal{O}(Jn^{3})+\mathcal{O}(\tilde{N}n^{2})$ operations for isotropic processes, and $\mathcal{O}(\tilde{N}n^{3})$ for anisotropic processes. As the likelihood function needs to be computed tens of times for numerically optimizing the parameters, the computational operations can be up to $10^{15}$ for a regular video with $500\times 500$ pixels and $500$ time frames, which is too computationally expensive. Here, we introduce a fast approach that can substantially reduce the computational cost without any approximation to the likelihood function.
@@ IV.2 Data reduction
For processes with an increased MSD with respect to the increase of lag time, such as subdiffusion or superdiffusion, the transformed intensity rapidly decorrelates at large wave vectors, making it indistinguishable from noise. A typical way for dimension reduction is to choose the first $J_{0}$ rings of Fourier-transformed intensity that explains the most variability of the data: $({\sum^{J_{0}}_{j=1}A_{j}})/({\sum^{J}_{j=1}A_{j}})\geq 1-\varepsilon$, where $\varepsilon$ is a small number, similar to the probabilistic principal component analysis [65]. For instance, choosing $\varepsilon=0.001$ means we have $99.9\%$ of the variability from the transformed intensity explained. As $A_{j}$ becomes close to 00 at large $j$ and the ‘ring’ gets larger for large $j$, such a choice can substantially reduce the storage and computational requirements by avoiding computing a large number of transformed intensities at high-frequency wave vectors. Another way is to ensure that we select a relatively large proportion ($J_{0}/J\geq\beta$) to retain enough information from the signal. For some confined processes, such as the Ornstein–Uhlenbeck (OU) process, where the MSD approaches a plateau at large lag time, using a large fraction of wave vectors is a safer choice, as the temporal correlation of transformed intensities does not decrease at large wave vectors. Having a more conservative threshold that selects a larger $J_{0}$, e.g. $\varepsilon=0.001$ and $\beta=0.5$, is also better for estimating the noise parameter. Here, we emphasize that the selection of $J_{0}$ is mainly for further reducing the storage and computational cost, which is different from selecting the range of wave vectors to analyze in minimizing the loss function in Eq. (4) in DDM. The default version of our packages utilize all wave vectors in estimation [33, 34].